OpenStudy (anonymous):

solve each equation on the interval [0,2pi] tan3x=square root 3

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

\[\tan(3x)=\sqrt{3}\]

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@perl

OpenStudy (sleepyhead314):



OpenStudy (anonymous):

it'll be pi/3 and pi/6?

OpenStudy (sleepyhead314):

to get (root(3)) / 1 which is what they want you have a tangent angle of pi/3 so 3x = pi/3

OpenStudy (anonymous):

okay i understand that part/

OpenStudy (anonymous):

what next?

OpenStudy (sleepyhead314):

what did you get for x?

OpenStudy (anonymous):

x=pi/9

OpenStudy (sleepyhead314):

then in what quadrants is tangent positive?

OpenStudy (anonymous):

3rd

OpenStudy (anonymous):

isnt there more than one equaiton